Halcyon Lines is considering the purchase of a new bulk carrier for $7.0 million. The forecasted revenues are $6.6 million a year and operating costs are $5.6 million. A major refit costing $3.6 million will be required after both the fifth and tenth years. After 15 years, the ship is expected to be sold for scrap at $3.1 million.

a. What is the NPV if the opportunity cost of capital is 10%?
Answer #1

The net profit for a year will be $6.6 million - $5.6 million = 1 million.

The NPV is calculated below:

NPV=\left [1\, million\times \left ( \frac{1-\left ( 1+0.10 \right )^{-15}}{0.10} \right )+\frac{3.1\, million}{1.10^{15}}-7\, million-\frac{3.6\, million}{1.10^{5}}-\frac{3.6\, million}{1.10^{10}} \right ]NPV=\left [1\, million\times 7.6060795+\frac{3.1\, million}{4.17724817}-7\, million-\frac{3.6\, million}{1.61051}-\frac{3.6\, million}{2.5937425} \right ]

NPV=\left [7.6060795\, million+0.7421153\, million-7\, million-2.2353167\, million-1.387956\, million ]NPV=-2.275078\, million